OpenStudy (anonymous):

a line passes through a point ( 4,1) and has a slope of 9/2 what is the equation of the line?

OpenStudy (anonymous):

y = 9/2x + b plug 4 = x and y = 1 in (coming from (4,1)) and solve

OpenStudy (anonymous):

would it be y=9/2x-17 or y=9/2-19

OpenStudy (anonymous):

the former, y=9/2x - 17